Integration of 6tan3x , limits [0 , 15] ?

Grade:11

3 Answers

Kinjal Shah
10 years ago
intgrl 6tna3x = 6((ln mod sec3x )/3) limits 0,15 =2( 1 + ln mod cos 45 )
Sunil Kumar FP
10 years ago
tan3x=-1/3ln cosx
therefore limit from0 to 15

-2lncos15

thanks and regards
sunil kr
askIItian faculty
Kinjal Shah
10 years ago
Sorry its: -2(1+ ln mod cos 45)